OpenStudy (anonymous):

The formula for the volume of a sphere is y = 4/3 pi r^3 . a. Solve the formula for r and write using positive rational exponents. b. Find the radius, to the nearest hundredth, of a sphere with a volume of 15 in.3

OpenStudy (tyteen4a03):

a. The idea is to make r the subject (r = something). The ^3 part in r needs to be gotten rid of too. b. Finish a, and plug in the numbers.

OpenStudy (anonymous):

can you help me with getting r by itself, im having trouble

OpenStudy (anonymous):

@tyteen4a03

OpenStudy (tyteen4a03):

y = 4/3 pi r^3 To start off, move everything that is not r to the LHS. This means we'll be dividing both sides by 4pi/3: 3y / 4pi = r^3. We know the inverse of r ^3 is r ^ 1/3 (i.e cube root), so cube root both sides and get the answer.

OpenStudy (anonymous):

I understood everything up to that point just not how to cube root 3y/pi

OpenStudy (tyteen4a03):

Basically wrap a cube root around it. \(\sqrt[3]{Stuff goes here}\)

OpenStudy (anonymous):

so I dont think this is right but I got y/1.5 = r

OpenStudy (tyteen4a03):

You should leave the surd roots intact. BTW, your answer is incorrect.

OpenStudy (anonymous):

oh okay. and you mean dont do anything to the left side?

OpenStudy (tyteen4a03):

@forty-two Basically if you see \(\sqrt[3]{xbr}\), leave this term as-is.